OpenStudy (anonymous):

given the equation y+10=- (x-15) what is the slope

OpenStudy (ilovebmth1234):

y= -x+15-10 y= -x +5 Putting it in the form of the equation(y=mx+or-c) We get: m= -1 Therefore slope= -1

OpenStudy (anonymous):

rearange the equation so that y=mx + b with m being the slope and b being the y intercept

OpenStudy (kc_kennylau):

That means trying to make y the subject

OpenStudy (ilovebmth1234):

yeah